Convert into mole.
     
       (c) 22 g of Carbon dioxide.
j,hgkmfthjrshbsxdn fjdgesbeased

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


Given mass of Carbon dioxide = 22 g
Molar mass of carbon dioxide (CO2) = (1 × 12) + (2 × 16)
= 12 + 32 = .44 g
Mole of carbon dioxide = 22/44 mole



Thanks